OpenStudy (anonymous):

Given the equation y − 3 =1/2 (x + 6) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (anonymous):

First: 1/2(x+6) = 1/2x + 3 Then: Add 3 to both sides ---> y = 1/2x + 6 Y = 1/2x + 6 is in standard form now!

OpenStudy (anonymous):

y − 3 =1/2 (x + 6) .....Distribute the 1/2 over y-3=(1/2)x+3 now add the 3 from the left side to the right side y=(1/2)x+6, now subtract the (1/2)x from the right side to the left side. -(1/2)x+y=6 multiply the whole equation by -1 .5x-y=-6..... This would be in standard form.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Yep